OpenStudy (anonymous):

solve 2(5^2x) - 7(5^x) = 15 for x. HELP!

OpenStudy (anonymous):

let 5^x=a then the equation will be 2a^2-7a=15 let all numbers in one side. 2a^2-7a-15=0. find a. then put 5^x=a find x.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome :) i wish it would be helpful.